OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 2x - y = 8. intercept = (2, 0); y-intercept = (0, -8) x-intercept = (4, 0); y-intercept = (0, -8) intercept = (2, 0); y-intercept = (0, 8) intercept = (4, 0); y-intercept = (0, 8)

OpenStudy (anonymous):

x-intercept :put y=0,2x-0=8,x=4,(4,0) y-x-intercept::put x=0,2(0)-y=8,y=-8,(0,-8) so the answer, x-intercept = (4, 0); y-intercept = (0, -8)

OpenStudy (anonymous):

THANK YOU!!!!!!! ={)

OpenStudy (anonymous):

^_________^